OpenStudy (anonymous):

Multiply. Try to recognize the type of product before multiplying (8a^8+5)(8a^8-5)

OpenStudy (anonymous):

(a+b)(a-b) looks similar, no? \[(a+b)(a-b) =a ^{2}-b ^{2}\]

OpenStudy (anonymous):

i try that in got myself lost

OpenStudy (anonymous):

a=8a^8 b=5 so just square them...

OpenStudy (anonymous):

8a^8(8a^8+5)+5(8a^8-5)

OpenStudy (anonymous):

a=8a^8 b=5 \[a ^{2} =( 8a ^{8})^{2} = 64a ^{16}\] \[b ^{2 } = 5^{2}=25\] \[a ^{2}-b ^{2}=64a ^{16}-25\]